Please explain how you would find the sum of the first 50 even numbers. What is the sum?
Answer by scott8148(6628)

the first 50 even numbers are 2 through 100

100+2=102 __ 98+4=102 __ 96+6=102 ... this pattern continues giving 25 sums of 102 each

25*102=2550